OpenStudy (anonymous):

For an amortized loan, you are given the loan, annual interest rate, and length of the loan. Find the monthly payment necessary to pay off the loan. Do not round until the final answer. Then, round to the nearest cent. amount: $9,000; rate: 9.6%; time: 3 years

jimthompson5910 (jim_thompson5910):

Use the formula P = L[(r/n)*(1 + r/n)^(n*t)]/[(1 + r/n)^(n*t) - 1] where L = amount loaned out (or amortized) r = annual interest rate n = compounding frequency t = # of years in this case L = 9000 r = 0.096 n = 12 (usually compounded monthly) t = 3

OpenStudy (anonymous):

Im not getting the right answer please help! jim_thompson5910

jimthompson5910 (jim_thompson5910):

P = L((r/n)*(1 + r/n)^(n*t))/((1 + r/n)^(n*t) - 1) P = 9000((0.096/12)*(1 + 0.096/12)^(12*3))/((1 + 0.096/12)^(12*3) - 1) P = 288.717440845243 P = 288.72

jimthompson5910 (jim_thompson5910):

So the monthly payment is $288.72

OpenStudy (anonymous):

can you help with this one too? For an amortized loan, you are given the loan, annual interest rate, and length of the loan. Find the monthly payment necessary to pay off the loan. Do not round until the final answer. Then, round to the nearest cent. Amount: $3,200; rate: 12%; time: 2 years.

jimthompson5910 (jim_thompson5910):

use the same formula but now L = 3200 r = 0.12 n = 12 t = 2

OpenStudy (anonymous):

is the answer 150.64??

jimthompson5910 (jim_thompson5910):

let me check

jimthompson5910 (jim_thompson5910):

P = L((r/n)*(1 + r/n)^(n*t))/((1 + r/n)^(n*t) - 1) P = 3200((0.12/12)*(1 + 0.12/12)^(12*2))/((1 + 0.12/12)^(12*2) - 1) P = 150.635111114447 P = 150.64 So you are correct

OpenStudy (anonymous):

what about this one? John bought a new computer for $1,650. He paid a $160 down payment and financed the rest for 1 year at an interest rate of 7%. Find the total interest paid on the given amortized loan assuming that John makes monthly payments. Do not round until the final answer. Then, round to the nearest cent.

jimthompson5910 (jim_thompson5910):

first find the monthly payment and tell me what that is

jimthompson5910 (jim_thompson5910):

keep in mind he's not financing all of the 1650

OpenStudy (anonymous):

im lost

jimthompson5910 (jim_thompson5910):

he's financing 1650 - 160 = ???

OpenStudy (anonymous):

1490

jimthompson5910 (jim_thompson5910):

so L = 1490

jimthompson5910 (jim_thompson5910):

r = 0.07

jimthompson5910 (jim_thompson5910):

n = 12 t = 1

jimthompson5910 (jim_thompson5910):

plug them all in

jimthompson5910 (jim_thompson5910):

then tell me what the monthly payment is

OpenStudy (anonymous):

57.10?

jimthompson5910 (jim_thompson5910):

that's too low, try again

OpenStudy (anonymous):

67.76?

jimthompson5910 (jim_thompson5910):

still too low, show me your work of what you're typing in

OpenStudy (anonymous):

the answer choices doesnt list anything higher than 67.76

jimthompson5910 (jim_thompson5910):

I don't care about the final answer right now

jimthompson5910 (jim_thompson5910):

I just want to know what the monthly payments are

jimthompson5910 (jim_thompson5910):

this might help if you're completely stuck http://www.bankrate.com/calculators/mortgages/loan-calculator.aspx it won't give you the final answer, but it will give you the monthly payment

jimthompson5910 (jim_thompson5910):

but I recommend using the formula I gave you and practice practice practice once you get the hang of things enough, then you can use the calculator

OpenStudy (anonymous):

is it 124.21??

jimthompson5910 (jim_thompson5910):

closer, but still a bit low

jimthompson5910 (jim_thompson5910):

did you use that calculator to confirm?

OpenStudy (anonymous):

what do i put for the loan start rate?

jimthompson5910 (jim_thompson5910):

leave that alone

jimthompson5910 (jim_thompson5910):

it doesn't matter when you start the loan date (so you can do it today or whenever)

jimthompson5910 (jim_thompson5910):

that's just used for setting up payment tables (assuming you actually took out a real loan)

OpenStudy (anonymous):

i plugged everything in but its still giving me 124.21

jimthompson5910 (jim_thompson5910):

see attached of what you were supposed to plug in (and what monthly payment you're supposed to get)

jimthompson5910 (jim_thompson5910):

do you see how to get that?

OpenStudy (anonymous):

ohh okay i put in .07 instead of 7

jimthompson5910 (jim_thompson5910):

i gotcha

jimthompson5910 (jim_thompson5910):

so you make monthly payments of 128.92 for 12 months to get (128.92)*(12) = 1547.04

jimthompson5910 (jim_thompson5910):

so you borrow 1490 dollars and pay back 1547.04 dollars total interest paid = 1547.04 - 1490 = 57.04

OpenStudy (anonymous):

wait so is 57.04 the answer?

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

thats not one of the choices?

OpenStudy (anonymous):

theres 57.10 though thats the closest

jimthompson5910 (jim_thompson5910):

what are all your choices?

OpenStudy (anonymous):

52.66 63.84 67.76 57.10

jimthompson5910 (jim_thompson5910):

hmm there has to be a typo then, idk

OpenStudy (anonymous):

okay thank you anyways

jimthompson5910 (jim_thompson5910):

np